The constant force F with components F_x = 3 N and
Answer #1

Work done is given by:

W = F.d

F = 3 i + 4 j

d = (14 - 2) i + (1 - 6) j

d = 12 i - 5 j

W = (3 i + 4 j).(12 i - 5 j)

W = 3*12 - 4*5 = 16 J

Correct option is A.